lilbene:

0.74 kilocalories per minute to calories per second ( Scientific Notation) I know 1000 is the calories and then it would be 740 so would it be 7.40 x 10 ^2??

mindfreeze:

you forgot to divide it by 60 to get minutes into seconds

lilbene:

would I divide 740 by 60 or before i moved the decimal?

mindfreeze:

you can do it anytime to be honest

AZ:

7.4 * 10^2 is the same as 740 but it's easier to divide 740 because you won't get confused with any of the decimals and you can figure it out properly after you divide it by 60

lilbene:

ok so id get 12.3333333 id move it once i believe to make it 1.23333333 I think ?

AZ:

and what's the exponent on 10 if you're moving it over once to the left

mindfreeze:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @lilbene ok so id get 12.3333333 id move it once i believe to make it 1.23333333 I think ? \(\color{#0cbb34}{\text{End of Quote}}\) 12.3333 is right

mindfreeze:

thank you for the medal

lilbene:

-1? @AZ

AZ:

Are you sure? \( 10^{-1} = \dfrac{1}{10^1} = \dfrac{1}{10}\) when you move the decimal to the left, the exponent should be positive like when you move it two times to the left in 740 to get 7.40 * 10^2 If you're moving the decimal to the RIGHT, then the exponent is going to be negative

lilbene:

oh ok your right so it is 1.23333333 x 10 ^ 1 ?

AZ:

You got it!

AZ:

Have you learned about significant figures?

lilbene:

yes I have just not much about the different conversions and its been a while i learned this around october!

AZ:

I'm just trying to see if you should cut down on some of those digits or if it doesn't matter when you submit that answer

AZ:

Ah okay, because we started off with 0.74 which has two significant figures so our final answer should also have 2 sig figs unless the teacher doesn't care then you could just put as many digits as you want haha

lilbene:

I honestly don't know she never mentioned anything about how many numbers she would like us to have but understandable a lot of 3"s !

AZ:

It's a never ending 3 that's why haha but you can definitely drop a few. From what I've noticed, they really only care about the sig figs when they teach that in the beginning of the year and then after that you just move on haha

AZ:

especially when you have multiple choice, it becomes irrelevant

lilbene:

your right, and I wish she doesn't believe in multiple choice which makes everything harder :(

lilbene:

I always 2nd guess myself

AZ:

You shouldn't! If you're able to calculate some number that's the same as the answer choice, you're most likely right! Unless you make like a common mistake which is what they're testing- to make sure you're able to use the formula correctly But if you know your stuff, you'll be fine!!